OpenStudy (anonymous):

−2x + 7y = 13 3x + 2y = −7

jhonyy9 (jhonyy9):

multiply first equation by 3 and second by 2 -6x +21y=39 6x +4y = -14 ------------- 0 25y =25 y=1 -2x+7(1)=13 -2x=6 x=6/-2 x=-3 x= -3 y= 1